# Towards the Causal Complete Cause of Multi-Modal Representation Learning

### v4

@@ Abstract
Multi-Modal Learning (MML) aims to learn effective representations across modalities for accurate predictions. Existing methods typically focus on modality consistency and specificity to learn effective representations. However, from a causal perspective, they may lead to representations that contain insufficient and unnecessary information. To address this, we propose that effective MML representations should be causally sufficient and necessary. Considering practical issues like spurious correlations and modality conflicts, we relax the exogeneity and monotonicity assumptions prevalent in prior works and explore the concepts specific to MML, i.e., Causal Complete Cause ($C^{3}$). We begin by defining $C^{3}$, which quantifies the probability of representations being causally sufficient and necessary. We then discuss the identifiability of $C^{3}$ and introduce an instrumental variable to support identifying $C^{3}$ with non-exogeneity and non-monotonicity. Building on this, we conduct the $C^{3}$ measurement, i.e., $C^{3}$ risk. We propose a twin network to estimate it through (i) the real-world branch: utilizing the instrumental variable for sufficiency, and (ii) the hypothetical-world branch: applying gradient-based counterfactual modeling for necessity. Theoretical analyses confirm its reliability. Based on these results, we propose $C^{3}$ Regularization, a plug-and-play method that enforces the causal completeness of the learned representations by minimizing $C^{3}$ risk. Extensive experiments demonstrate its effectiveness.
@@ 1 Introduction
The initial inspiration for artificial intelligence is to imitate human perceptions which are based on different modalities (Jordan & Mitchell, 2015; Mahesh, 2020), e.g., sight, sound, and touch. Each modality provides unique information with distinct properties (Baltrušaitis et al., 2018; Wang et al., 2023c). Multi-modal learning (MML) (Huang et al., 2021a; Ge et al., 2023) has emerged as a promising approach to emulate human perception. It aims to learn good representations from multiple modalities for accurate decisions.
In this context, a key question arises: what defines a “good” MML representation? Existing methods typically define good MML representations from two perspectives: modality consistency and modality specificity (Zhang et al., 2023a; Ge et al., 2023; Radford et al., 2021; Fan et al., 2024; Dong et al., 2024). The former emphasizes extracting modality-shared semantics related to primary events within the MML task. This type of methods (Liang et al., 2022; Xia et al., 2024; Abdollahzadeh et al., 2021) aims to obtain unified representations by mapping features from different modalities into a common embedding space. In contrast, the second perspective believes modality specificity captures the distinct statistical properties of each modality, reflecting different aspects of the primary events. This type of methods (Dong et al., 2024; Yang et al., 2022; Frost et al., 2015; Zhou et al., 2021) decomposes features within each modality into modality-specific and modality-shared components, learning all shared components while applying distance constraints on modality-specific features to enhance diversity.
@@
However, from a causal perspective (Ahuja et al., 2020; Koyama & Yamaguchi, 2020), these methods may result in the learned representations being insufficient or unnecessary. Specifically, sufficiency indicates that the use of the representations will establish the label, while necessity indicates that the label becomes incorrect when the representations are absent (Pearl, 2009). Figure 1 provides an example. If the MML model only focuses on causal sufficiency, it will lose important modality-specific semantics, affecting generalization; if the model only focuses on causal necessity, the decisions may be made incorrectly based on the background, affecting discriminability. Correspondingly, considering the aforementioned two types of MML methods: (i) those focused on modality consistency may cause the model to overlook important modality-specific semantics; (ii) those focused on modality specificity may incorrectly capture irrelevant information. Thus, existing MML methods (Dong et al., 2024; Lu et al., 2022) without causal constraints may fail to satisfy sufficiency and necessity, affecting model performance. The experiments in Section 6.2 further prove this (Figure 4 and Table 1): (i) the representations learned by existing methods have much lower correlation scores with sufficient and necessary causes than the proposed method with specific constraints; (ii) after constraining the learned representations, the performance of existing MML methods are significantly improved. The analyses in Section 4 also emphasize the importance of causal sufficiency and necessity. Thus, inspired by (Pearl, 2009; Yang et al., 2024), we propose that a good MML representation must with both causal sufficiency and necessity, i.e., being causal complete.
@@
Noticeably, in practical MML applications, constraining the causal sufficiency and necessity of learned representations remains challenging. Specifically, the related discussion in existing works (Yang et al., 2024; Pearl, 2009) are based on the exogeneity and monotonicity assumptions (Definition D.1 and Definition D.2 in Appendix D.4). Exogeneity refers to the scenario where the influence of the external intervention on the conditional distributions is negligible when the causal representation variable is exogenous relative to label variable, while monotonicity illustrates the consistent, unidirectional effect on the label variable of representation variable. However, in practice, non-trivial corner cases often arise: the inseparability of MML semantics leads to spurious correlations (Figure 2), and cross-modal conflicts combined with high-dimensional nonlinear interactions (Huang et al., 2021b; Wang et al., 2023d) undermine monotonicity. These challenges render the conditions of traditional causal sufficiency and necessity discussion inapplicable.
@@
To address this, in this paper, we relax the above assumptions to explore the causal sufficiency and necessity in MML, ensuring the quality of the learned representation. Firstly, we propose the definition of causal sufficiency and necessity, i.e., causal complete causes ($C^{3}$), for MML based on (Pearl, 2009). It reflects the probability that the learned representation is causally complete by estimating the probability of label change after an intervention on the representation given two conditions of observations. One condition is for sufficiency, and another is for necessity. Then, we analyze the identifiability of $C^{3}$, which allows us to quantify $C^{3}$ using the observable data in practice. Unlike previous studies, we propose an instrumental variable to ensure estimating the segmented effect of $C^{3}$ without spurious correlations, thereby relaxing the assumptions of exogeneity and monotonicity. Based on this, we propose the measurement of $C^{3}$ using the twin network, i.e., $C^{3}$ risk, where a low $C^{3}$ risk means that the learned representations are causally complete with high confidence. The challenge of $C^{3}$ measurement lies in eliminating the spurious correlations in the sufficiency evaluation and modeling the counterfactual data required for the necessity evaluation. The proposed twin network addresses this through (i) the real-world branch: using the proposed instrumental variable to eliminate spurious correlations in the representation, and (ii) the hypothetical-world branch: using provable gradient-based adjustments to model the counterfactuals. Through theoretical analyses, we prove the reliability of the twin network and provide the performance guarantee of the $C^{3}$ risk. Based on these theoretical results, we propose Causal Complete Cause Regularization ($C^{3}$), a plug-and-play method to learn causal complete representations by constraining their $C^{3}$ risks.
@@
The main contributions are as follows: (i) We propose the definition, identifiability, and measurement of the causal complete cause ($C^{3}$) concept for MML without the assumptions of exogeneity and monotonicity. These results provide an effective way, i.e., $C^{3}$ risk, to estimate the sufficiency and necessity of the learned representation for robust and accurate MML (Section 3). (ii) We theoretically demonstrate the effectiveness and reliability of the proposed $C^{3}$ risk. Inspired by this, we propose $C^{3}$, which can be applied to any MML model to learn causal complete representations with low $C^{3}$ risk ((Sections 4 and 5). (iii) Finally, we conduct extensive experiments on various datasets that prove the effectiveness and robustness of $C^{3}$ (Section 6).
@@ 2 Problem Analysis
Let $\mathcal{X}$, $\mathcal{Y}$, and $\mathcal{Z}$ represent the input space, label space, and latent space, respectively. Consider a MML task involving data $(x,y)\in\mathcal{X}\times\mathcal{Y}$, where $x\in\mathcal{X}$ denotes the sample and $y\in\mathcal{Y}$ denotes the label. All the data are sampled from the same joint distribution $P_{XY}$ over $\mathcal{X}\times\mathcal{Y}$. The training dataset is defined as $\mathcal{D}_{tr}=\{(x_{i},y_{i})\}_{i=1}^{N}$, and the testing dataset is defined as $\mathcal{D}_{te}$. Here, each sample $x_{i}=\{x^{(1)}_{i},\dots,x^{(K)}_{i}\}$ contains $K$ modalities. The objective of MML is to obtain a robust model $f_{\theta}=\mathcal{W}\circ g$ that performs well on unseen test dataset $\mathcal{D}_{te}$. Here, the $g:\mathcal{X}\rightarrow\mathcal{Z}$ is the feature extractor and $\mathcal{W}:\mathcal{Z}\rightarrow\mathcal{Y}$ is the classifier. The specific implementation is illustrated in Appendix G, and the table of notations is shown in Table 4.
From the perspective of data generation (Suter et al., 2019; Hu et al., 2022b), each sample is composed of various generating factors, e.g., color, shape, etc. In the context of MML, each sample contains generating factors from all modalities. From a causal perspective (Pearl, 2009), these generating factors can be divided into task-related factors that come from the label, i.e., causal factors $\text{F}_{c}$, and task-independent factors, i.e., non-causal factors $\text{F}_{s}$. The key to accurate prediction lies in extracting the causal generating factors $\text{F}_{c}$ that are related to labels in the latent space (Deshpande et al., 2022). Thus, the goal of $f_{\theta}$ can be re-established as learning representations that contain all causal generating factors $\text{F}_{c}$ from the data without $\text{F}_{s}$ for accurate prediction.
@@
As shown in Figure 1, assume that the training data is with three modalities of image, text, and audio, and all the modalities have the feature “duck paws” when a “duck” exists. The goal of this MML task is to classify the “duck”. Then, the models that learn representations based on consistency will capture the features of “duck paws”, and can establish the label “duck” based on the learned representations. However, the learned model may also make errors in another “duck” scenario where the MML samples do not contain “duck paws”, e.g., “duck swim on the lake” (upper right in Figure 1). This suggests that the learned representation contains sufficient but unnecessary causes since the label “duck” can be predicted using the current representation, but may not work in another scenario. Similarly, there are some necessary but insufficient representations, e.g., “duck” must with “wings”, but “wings” may also correspond to another label, e.g., “bird” (lower of Figure 1). For sufficient and necessary causes, they ensure that the learned representations not only reflect the features target the “duck” label, e.g., “duck paws”, but also the features where the “duck” label must have, e.g., “wings”. Thus, good representations must have both causal sufficiency and necessity, i.e., causal complete causes, which is the goal we explore in this study.
@@
We conduct a Structural Causal Model (SCM) for MML based on causal generating mechanism  (Suter et al., 2019; Deshpande et al., 2022), as shown in Figure 2 left. In this SCM, Y and X denote the label variable and corresponding generated data variable in the MML task. $\text{F}_{c}$ and $\text{F}_{s}$ represent the distinct sets of generating factors that are causally and non-causally related to Y. Each generating factor corresponds to a semantic of the MML task, e.g., color, shape, background, modality indicator, etc. Since both $\text{F}_{c}$ and $\text{F}_{s}$ represent high-level knowledge of the data but only $\text{F}_{c}$ is causally related to Y, we could naturally define the MML task label variable Y as the cause of the $\text{F}_{c}$, i.e., ${\text{Y}}\rightarrow\text{F}_{c}$. There is no connection between $\text{F}_{s}$ and Y. Following (Hu et al., 2022a; Deshpande et al., 2022), the sample in each modality is generated simultaneously using all the generating factors, including factors caused from the label and unobservable variables, e.g., environmental effects. The unobservable variable results in the generating factors $\text{F}_{s}$, which is non-causally related to Y. Then, we get both $\text{F}_{c}\rightarrow\text{X}$ and $\text{F}_{s}\rightarrow\text{X}$. Thus, we obtain Figure 2 left. Based on this, we further construct a SCM to discuss the learning process of MML, as shown in Figure 2 right. It can be viewed as the inverse process of the causal generating mechanism. Based on Figure 2 left, an ideal MML model should only utilize causal generating factors $\text{F}_{c}$ and be invariant to any intervention on non-causal generating factors $\text{F}_{s}$. However, in practice, $\text{F}_{c}$ and $\text{F}_{s}$ in the learned representation may be coupled (Dong et al., 2024; Li et al., 2023). This leads to the model potentially learning based on non-causal $\text{F}_{s}$. There exist spurious correlations between $\text{F}_{s}$ and Y, i.e., we get the additional $\text{F}_{s}\rightarrow\text{Y}$ in Figure 2 right.
The presence of spurious correlations leads to inaccuracies in MML. Specifically, the learned MML representations may fall into four scenarios: (i) sufficient and necessary: contains all $\text{F}_{c}$; (ii) necessary but insufficient: contains part of $\text{F}_{c}$; (iii) sufficient but unnecessary: contains all $\text{F}_{c}$ with $\text{F}_{s}$; (iv) insufficient and unnecessary: only contains $\text{F}_{s}$. See Appendix D and G for more analyses. To ensure accurate and robust MML, in this paper, we aim to propose a methodology that constrains the model to learn representations only contain causally sufficient and necessary information.

### v6

@@ Abstract
Multi-Modal Learning (MML) aims to learn effective representations across modalities for accurate predictions. Existing methods typically focus on modality consistency and specificity to learn effective representations. However, from a causal perspective, they may lead to representations that contain insufficient and unnecessary information. To address this, we propose that effective MML representations should be causally sufficient and necessary. Considering practical issues like spurious correlations and modality conflicts, we relax the exogeneity and monotonicity assumptions prevalent in prior works and explore the concepts specific to MML, i.e., Causal Complete Cause ($C^{3}$). We begin by defining $C^{3}$, which quantifies the probability of representations being causally sufficient and necessary. We then discuss the causal identifiability of $C^{3}$ and introduce an instrumental variable to support identifying $C^{3}$ with non-exogeneity and non-monotonicity. Building on this, we conduct the $C^{3}$ measurement, i.e., $C^{3}$ risk. We propose a twin network to estimate it through (i) the real-world branch: utilizing the instrumental variable for sufficiency, and (ii) the hypothetical-world branch: applying gradient-based counterfactual modeling for necessity. Theoretical analyses confirm its reliability. Based on these results, we propose $C^{3}$ Regularization, a plug-and-play method that enforces the causal completeness of the learned representations by minimizing $C^{3}$ risk. Extensive experiments demonstrate its effectiveness.
@@ 1 Introduction
The initial inspiration for artificial intelligence is to imitate human perceptions which are based on different modalities (Jordan & Mitchell, 2015; Mahesh, 2020), e.g., sight, sound, and touch. Each modality provides unique information with distinct statistical properties (Baltrušaitis et al., 2018; Wang et al., 2023c). A fundamental aspect of human perception is the ability to simultaneously integrate data from different modalities to understand the world (Xu et al., 2023). Multi-modal learning (MML) (Huang et al., 2021a; Ge et al., 2023) has emerged as a promising approach to emulate human sensory perception. It aims to learn good representations from multiple modalities that can achieve accurate decisions.
In this context, a key question arises: what defines a “good” MML representation? Existing methods typically define good MML representations from two perspectives: modality consistency and modality specificity (Zhang et al., 2023a; Ge et al., 2023; Radford et al., 2021; Fan et al., 2024; Dong et al., 2024). The former emphasizes extracting modality-shared semantics related to primary events within the MML task. This type of method (Liang et al., 2022a; Xia et al., 2024; Abdollahzadeh et al., 2021) aims to obtain unified representations by mapping features from different modalities into a common embedding space. In contrast, the second perspective believes modality specificity captures the distinct statistical properties of each modality, reflecting different aspects of the primary events. This type of method (Dong et al., 2024; Yang et al., 2022; Frost et al., 2015; Zhou et al., 2021) decomposes features within each modality into modality-specific and modality-shared components, learning all shared components while applying distance constraints on modality-specific features to enhance diversity.
@@
However, from a causal perspective (Ahuja et al., 2020; Koyama & Yamaguchi, 2020), these methods may result in the learned representations being insufficient or unnecessary. Specifically, sufficiency indicates that the use of the representations will establish the label, while necessity indicates that the label becomes incorrect when the representations are absent (Pearl, 2009). Figure 1 provides an example. If the MML model only focuses on causal sufficiency, it will lose important modality-specific semantics, affecting generalization; if the model only focuses on causal necessity, the decisions may be made incorrectly based on the background, affecting discriminability. Correspondingly, considering the aforementioned two types of MML methods: (i) those focused on modality consistency may cause the model to overlook important modality-specific semantics; (ii) those focused on modality specificity may incorrectly capture irrelevant information. Thus, existing MML methods (Dong et al., 2024; Lu et al., 2022) without causal constraints may fail to satisfy sufficiency and necessity, affecting model performance. The experiments in Section 6.2 further prove this (Figure 4 and Table 1): (i) the representations learned by existing methods have much lower correlation scores with sufficient and necessary causes than the proposed method with causal constraints; (ii) after constraining the learned representations, the performance of existing methods are significantly improved. The analyses in Section 4 also emphasize the importance of causal sufficiency and necessity. Thus, inspired by (Pearl, 2009; Yang et al., 2024), we propose that a good MML representation must be both causally sufficient and necessary, i.e., causally complete.
@@
Noticeably, in practical MML applications, constraining the causal sufficiency and necessity of learned representations remains challenging. Specifically, the related discussion in existing works (Yang et al., 2024; Pearl, 2009) is based on the exogeneity and monotonicity assumptions (Definition D.1 and Definition D.2 in Appendix D.3). Exogeneity refers to the scenario where the influence of the external intervention on the conditional distributions is negligible when the causal representation variable is exogenous relative to the label variable, while monotonicity illustrates the consistent, unidirectional effect on the label of the representation. However, in practice, non-trivial corner cases often arise: the inseparability of MML semantics leads to spurious correlations (Figure 2), and cross-modal conflicts combined with high-dimensional nonlinear interactions (Huang et al., 2021b; Wang et al., 2024) undermine monotonicity. These challenges render the conditions of traditional causal sufficiency and necessity discussion inapplicable.
@@
To address this, in this paper, we relax the above assumptions to explore the causal sufficiency and necessity in MML, ensuring the quality of the learned representation. Firstly, we propose the definition of causal sufficiency and necessity, i.e., causal complete causes ($C^{3}$), for MML based on (Pearl, 2009). It reflects the probability that the learned representation is causally complete by estimating the probability of label change after an intervention on the representation, given two conditions of observations. One condition is for sufficiency, and another is for necessity. Then, we analyze the causal identifiability of $C^{3}$, which allows us to quantify $C^{3}$ using the observable data in practice. Unlike previous studies, we propose an instrumental variable to ensure estimating the segmented effect of $C^{3}$ without spurious correlations, thereby relaxing the assumptions of exogeneity and monotonicity. Based on this, we propose the measurement of $C^{3}$ using the twin network, i.e., $C^{3}$ risk, where a low $C^{3}$ risk means that the learned representations are causally complete with high confidence. The challenge of $C^{3}$ measurement lies in eliminating the spurious correlations in the sufficiency evaluation and modeling the counterfactual data required for the necessity evaluation. The proposed twin network addresses this through (i) the real-world branch: using the proposed instrumental variable to eliminate spurious correlations in the representation, and (ii) the hypothetical-world branch: using provable gradient-based adjustments to model the counterfactuals. Through theoretical analyses, we prove the reliability of the twin network and provide the performance guarantee of the $C^{3}$ risk. Based on these theoretical results, we propose Causal Complete Cause Regularization ($C^{3}$), a plug-and-play method to learn causal complete representations by constraining their $C^{3}$ risks.
@@
The main contributions are as follows: (i) We propose the definition, identifiability, and measurement of the causal complete cause ($C^{3}$) concept for MML without the assumptions of exogeneity and monotonicity. These results provide an effective way, i.e., $C^{3}$ risk, to estimate the sufficiency and necessity of the learned representation for robust and accurate MML (Section 3). (ii) We theoretically demonstrate the effectiveness and reliability of the proposed $C^{3}$ risk. Inspired by this, we propose $C^{3}$, which can be applied to any MML model to learn causal complete representations with low $C^{3}$ risk (Sections 4 and 5). (iii) Finally, we conduct extensive experiments on various datasets that prove the effectiveness and robustness of $C^{3}$ (Section 6).
@@ 2 Problem Analysis
Let $\mathcal{X}$, $\mathcal{Y}$, and $\mathcal{Z}$ represent the input space, label space, and latent space, respectively. Consider a MML task involving data $(x,y)\in\mathcal{X}\times\mathcal{Y}$, where $x\in\mathcal{X}$ denotes the sample and $y\in\mathcal{Y}$ denotes the label. All the data are sampled from the same joint distribution $P_{XY}$ over $\mathcal{X}\times\mathcal{Y}$. The training dataset is defined as $\mathcal{D}_{tr}=\{(x_{i},y_{i})\}_{i=1}^{N}$, and the testing dataset is defined as $\mathcal{D}_{te}$. Here, each sample $x_{i}=\{x^{(1)}_{i},\dots,x^{(K)}_{i}\}$ contains $K$ modalities. The objective of MML is to obtain a robust model $f_{\theta}=\mathcal{W}\circ g$ that performs well on the unseen test dataset $\mathcal{D}_{te}$. Here, the $g:\mathcal{X}\rightarrow\mathcal{Z}$ is the feature extractor and $\mathcal{W}:\mathcal{Z}\rightarrow\mathcal{Y}$ is the classifier. The specific implementation is illustrated in Appendix G, and the table of notations is shown in Table 4.
From the perspective of data generation (Suter et al., 2019; Hu et al., 2022b), each sample is considered to be generated using a set of generating factors, e.g., color, shape, etc. In the context of MML, each sample contains generating factors from all modalities. From a causal perspective (Pearl, 2009), these generating factors can be divided into task-related factors, i.e., causal factors $\text{F}_{c}$, and task-independent factors, i.e., non-causal factors $\text{F}_{s}$. Among them, the causal factors are related to the task label, supporting accurate predictions (Deshpande et al., 2022). Thus, the goal of $f_{\theta}$ can be re-established as learning representations that contain all causal generating factors $\text{F}_{c}$ from the data without $\text{F}_{s}$.
@@
As shown in Figure 1, assume the goal of this MML task is to classify the “duck”, the analysis is conducted under specific task and data conditions. The training data of this task contains three modalities, i.e., image, text, and audio, and all the modalities have the feature “duck paws”. Then, the models that learn representations based on consistency may capture the features of “duck paws”, and can establish the label “duck” based on “duck paws”. However, the model may make errors in another “duck” scenario where the MML samples do not contain “duck paws”, e.g., “duck swims on the lake” (upper right in Figure 1). This suggests that the learned representation contains sufficient but unnecessary causes: the label “duck” can be predicted using the current representation (Line 1 Left), but may not work in another scenario (Line 1 Right). Similarly, the models may also learn necessary but insufficient representations, e.g., the representations with feature “wings”. Take the Line 2 of Figure 1 as an example, “duck” must have “wings”, but the samples with “wings” may also correspond to another label, e.g., “bird” (lower of Figure 1), leading to incorrect predictions. Under this context, a sufficient and necessary feature can be “flat duck bill” as its presence indicates “duck” and every “duck” sample includes it. Briefly, for sufficient and necessary causes, they ensure that the learned representations not only reflect the features targeted the “duck” label, but also the features that the “duck” label must have. Thus, a good representation must have both causal sufficiency and necessity, i.e., causal complete causes, for accurate and robust prediction. This is also the goal we explore in this study. More analyses are provided in Appendix D.2.
@@
We construct a Structural Causal Model (SCM) for MML based on the causal generating mechanism  (Suter et al., 2019; Deshpande et al., 2022) in Figure 2 left. In this SCM, Y and X denote the label variable and corresponding generated data variable in the MML task. $\text{F}_{c}$ and $\text{F}_{s}$ represent the distinct sets of generating factors that are causally and non-causally related to Y. Each generating factor corresponds to a semantic of the MML task, e.g., color, shape, background, modality indicator, etc. Since both $\text{F}_{c}$ and $\text{F}_{s}$ represent high-level knowledge of the data but only $\text{F}_{c}$ is causally related to Y, we could naturally define the MML task label variable Y as the cause of the $\text{F}_{c}$, i.e., ${\text{Y}}\rightarrow\text{F}_{c}$. There is no connection between $\text{F}_{s}$ and Y. Following (Hu et al., 2022a; Deshpande et al., 2022), the sample in each modality is generated simultaneously using all the generating factors, including factors caused from the label and unobservable variables, e.g., environmental effects. The unobservable variable results in the generating factors $\text{F}_{s}$, which are non-causally related to Y. Then, we get both $\text{F}_{c}\rightarrow\text{X}$ and $\text{F}_{s}\rightarrow\text{X}$. Thus, we obtain Figure 2 left. Based on this, we further construct an SCM to discuss the learning process of MML, as shown in Figure 2 right. It can be viewed as the inverse process of the causal generating mechanism. Based on Figure 2 left, an ideal MML model should only utilize causal generating factors $\text{F}_{c}$ and be invariant to any intervention on non-causal generating factors $\text{F}_{s}$. However, in practice, $\text{F}_{c}$ and $\text{F}_{s}$ in the learned representation may be coupled (Dong et al., 2024; Li et al., 2023). This leads to the model potentially learning based on non-causal $\text{F}_{s}$. There exist spurious correlations between $\text{F}_{s}$ and Y, i.e., we get the additional $\text{F}_{s}\rightarrow\text{Y}$ in Figure 2 right.
The presence of spurious correlations leads to inaccuracies in MML. Specifically, the learned MML representations may fall into four scenarios: (i) sufficient and necessary: contains all $\text{F}_{c}$; (ii) necessary but insufficient: contains part of $\text{F}_{c}$; (iii) sufficient but unnecessary: contains all $\text{F}_{c}$ with $\text{F}_{s}$; (iv) insufficient and unnecessary: only contains $\text{F}_{s}$. See Appendix D and G for more analyses. To ensure accurate and robust MML, in this paper, we aim to propose a methodology that constrains the model to learn representations that contain only causally sufficient and necessary information.